OpenStudy (anonymous):

A recreation room has the following dimensions. Carpeting will cost $9.00 a square yard. How many square yards of carpet will be needed and how much will it cost? Include correct units at each stage within your solution. 3yd ,2yd ,2yd 4yd

OpenStudy (anonymous):

I don't understand the room dimensions.

OpenStudy (anonymous):

? dont understand

OpenStudy (anonymous):

is there a diagram of the room? is it a trapezoid? you have not given enough information to tell how to calculate the area of the room.

OpenStudy (anonymous):

what is the room area?

OpenStudy (anonymous):

rectangle triangle and a square all together

OpenStudy (anonymous):

ractangle values are 3 and 2 and square 4 and 3

OpenStudy (anonymous):

its a trapazoid hight^2 = e1^2 -( (b1-b2)/2)^2 h^2= 2^2 - .5^2 h^2= 4 - .25 h = \[\sqrt{3.25}\] A = h *(b1 + b2)/2 \[A =\sqrt{3.25} * (4 +3)/2\] \[A =3.5\sqrt{3.25}\]

OpenStudy (anonymous):

rockstar, first of all, 4 - .25 = 3.75. second of all, how do you know the room is a trapezoid?

OpenStudy (anonymous):

sry its like 3 am and im calculating in my head and trapazoid is most logical

OpenStudy (anonymous):

a trapezoidal rec room is most logical? I think we need a diagram. We're just guessing unless shenita gives us a better description or a picture.

OpenStudy (anonymous):

fromme the dimmensins given it is

OpenStudy (anonymous):

i cnt post pic but its those three in one figure made together

OpenStudy (anonymous):

you're not giving us enough information to determine the area of the room.